OpenStudy (anonymous):

Solve for x using factoring and zero product property: y=x^2-x-12 PLEASE HELP!

OpenStudy (anonymous):

solve for x ??? as in isolate the x in this equation???

OpenStudy (anonymous):

or do you mean to find the zeros of \(\large y=x^2-x-12 \) ???

OpenStudy (anonymous):

yes the second one

OpenStudy (anonymous):

ok... so we need to solve: \(\large 0 = x^2-x-12 \)

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

notice that: \(\large x^2-x-12=(x-4)(x+3) \) so the equation becomes: \(\large 0=(x-4)(x+3) \) isn't it obvious what x needs to be for this equation to hold true?

OpenStudy (anonymous):

ummm yes?

OpenStudy (anonymous):

so what should x = ???

OpenStudy (anonymous):

7??

OpenStudy (anonymous):

let's see... x=7 \(\large 0=(\color {red}7-4)(\color {red}7+3) \) is this a true statement?

OpenStudy (anonymous):

I have no idea anymore, I'm confused ...

OpenStudy (anonymous):

\(\large 0=(\color {red}7-4)(\color {red}7+3) \) \(\large 0=(3)(10) \) \(\large 0=30 \) is this a true statement?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

zero is never equal to 30... the answer is no.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Either one of the factors must be 0.